Consider the following competing hypotheses and accompanying sample data. (You may find it useful to reference the appropriate table: z table or t table) H0: p1 − p2 ≥ 0 HA: p1 − p2 < 0 x1 = 248 x2 = 266 n1 = 444 n2 = 444

What is the Test Statistic?
Answer #1

p1cap = X1/N1 = 248/444 = 0.5586
p1cap = X2/N2 = 266/444 = 0.5991
pcap = (X1 + X2)/(N1 + N2) = (248+266)/(444+444) = 0.5788

Test statistic
z = (p1cap - p2cap)/sqrt(pcap * (1-pcap) * (1/N1 + 1/N2))
z = (0.5586-0.5991)/sqrt(0.5788*(1-0.5788)*(1/444 + 1/444))
z = -1.22